To start, you will need to pick the CFFA for a subject firm.  Specifically, pick such that $3,000,000 < CFFA < $5,000,000. For this question, you will form an estimate of value for the following privately-held firm that is up for sale. Assume:

-The firm’s financing structure is 100% debt.  The firm’s cost of debt is 12% annually and it is not assumable.

-The firm’s CFFAgrow at a constant annual rate of 3%.

-Like-kind firms are trading at $50,000,000 with CFFAranging between $3,000,000 and $5,000,000 with constant growth of 4% in their CFFA.

-There is strong demand for firms of this type in the marketplace, i.e. assume the highest and best use principal applies.
Answer #1

What is the Question ??

If you have any CFFA value say x

PV (Perpetuity with 3% growth rate & 12% cost ) = x /(0.12-0.03) = x/0.09

say x = $ 4,000,000

Value = 4,000,000 /0.09 = 44,444,444.44

To have value of 50,000,000 ===> x = 50,000,000*0.09 = $4,500,000

x satisfies the condition that 3,000,000 < x < 5,000,000